OpenStudy (anonymous):

Write the equation of the line with a slope of 3/7 passing through the point (9, -2). a.2x-9y= -3/7 b.Y=3/7x+9 c.Y=3/7x-2 d. 3x - 7y = 41

OpenStudy (ghazi):

y+2= 3/7*(x-9)

OpenStudy (ghazi):

D

OpenStudy (jim766):

You could also use y = mx + b -2 = 3/7(9) + b -2 = 27/7 + b -27/7 -27/7 -41/7 = b y = 3/7x - 41/7 D. 3x - 7y = 41 -3x -3x -7y = -3x + 41 ___ ___ ___ -7 -7 -7 y = 3/7x - 41/7

OpenStudy (anonymous):

thanks